# Low-Rank Tensor Decomposition over Finite Fields

### v2

@@ Abstract
We show that finding rank-1, rank-2, and rank-3 decompositions of a 3D tensor over a fixed finite field can be done in polynomial time. However, if some cells in the tensor are allowed to have arbitrary values, then rank-2 is NP-hard over the integers modulo 2. We also explore rank-1 decomposition of a 3D tensor and of a matrix where some cells are allowed to have arbitrary values.
@@ 1.1 Results
For $R\leq 3$, finding a rank-$R$ decomposition over a finite field $\mathbb{F}$ of an $n\times n\times n$ tensor, or determining that it does not exist, can be done in $O(f(|\mathbb{F}|,R)n^{3})$ time, for $f$ polynomial in $|\mathbb{F}|$ and exponential in $R$.
@@ 2 Without wildcards
In this section we describe our algorithm for Theorem 1. The following observations of rank-$\leq 2$ decomposition over $\mathbb{Z}/2\mathbb{Z}$ serve as inspiration:
  • Rank-1 is easy since $abc=1$ forces $a=1$, $b=1$, and $c=1$, so $T=a\times b\times c$ if and only if $(a_{i}=\lor_{j,k}T_{i,j,k},\ b_{j}=\lor_{i,k}T_{i,j,k},\ c_{k}=\lor_{i,j}T_{i,j,k})$ is a solution.
  • Rank-2 at first glance is not so clear, but we notice that if $T=a_{0}\times b_{0}\times c_{0}+a_{1}\times b_{1}\times c_{1}$, then all slices $T_{i,:,:}$ are of the form $u(b_{0}\times c_{0})+v(b_{1}\times c_{1})$ for $u,v\in\mathbb{Z}/2\mathbb{Z}$, so each $T_{i,:,:}$ can only be one of four possibilities. These slices also have rank at most 2, so one could maybe do exhaustive casework over the ranks of these slices to find a rank-2 decomposition of $T$ (or prove that one does not exist).
Our algorithm generalizes to rank-3 and arbitrary finite fields. Suppose we want to find a rank-$R$ decomposition of the $n\times n\times n$ tensor $T$ over the finite field $\mathbb{F}$, i.e., $T_{i,j,k}=\sum_{r=0}^{R-1}A_{r,i}B_{r,j}C_{r,k}\ \forall i,j,k$. Our main insight is to fix every possible $A$ and then solve for $B,C$ in a clever manner. To ensure we do not enumerate too many $A$ (specifically, the number of possible $A$ we search over is $O(1)$ w.r.t. $n$), we extract a basis on the slices $T_{i,:,:}$.
Flatten each $T_{i,:,:}$ to form the $n\times n^{2}$ matrix $\overline{T}:=\left[T_{i,:,:}\texttt{.flatten}()\right]_{i}$111 We use numpy.ndarray.flatten to flatten an $n\times n$ matrix into a $n^{2}$-length vector, with elements in row-major order. For example, $\begin{bmatrix}a&b\\ c&d\end{bmatrix}\texttt{.flatten}()=\begin{bmatrix}a&b&c&d\end{bmatrix}$. ; run Gaussian elimination to get $\overline{T}=M\mathrm{rref}\left(\overline{T}\right)$, where rref denotes reduced row-echelon form and $M$ is invertible; let $J$ be the number of nonzero rows of $\mathrm{rref}\left(\overline{T}\right)$, which is equal to $\mathrm{rank}\left(\overline{T}\right)$. Then we can collect basis slices to form the tensor $T^{\prime}=\left[\mathrm{rref}\left(\overline{T}\right)_{j,:}\texttt{.reshape}(n,n)\right]_{0\leq j<J}$222 We use numpy.ndarray.reshape to reshape a $n^{2}$-length vector into a $n\times n$ matrix, with elements in row-major order. For example, $\begin{bmatrix}a&b&c&d\end{bmatrix}\texttt{.reshape}(2,2)=\begin{bmatrix}a&b\\ c&d\end{bmatrix}$. , yielding $T_{i,:,:}=\sum_{0\leq j<J}M_{i,j}T^{\prime}_{j,:,:}$ and $T^{\prime}_{j,:,:}=\sum_{0\leq i<n}(M^{-1})_{j,i}T_{i,:,:}$. These relations allow us to convert between decompositions of $T$ and $T^{\prime}$:
$T^{\prime}=\sum_{r=0}^{R-1}U_{r}\times B_{r,:}\times C_{r,:}$ $\Rightarrow T=\sum_{r=0}^{R-1}\left[\sum_{0\leq j<J}M_{i,j}U_{r,j}\right]_{i}\times B_{r,:}\times C_{r,:}$
$T=\sum_{r=0}^{R-1}A_{r}\times B_{r,:}\times C_{r,:}$ $\Rightarrow T^{\prime}=\sum_{r=0}^{R-1}\left[\sum_{0\leq i<n}(M^{-1})_{j,i}A_{r,i}\right]_{i}\times B_{r,:}\times C_{r,:}$
Thus, finding a rank-$R$ decomposition of $T$ is equivalent to finding a rank-$R$ decomposition of $T^{\prime}$. We can then rewrite $T^{\prime}=\sum_{r}U_{r}\times B_{r}\times C_{r}$ into the block matrix form $\begin{bmatrix}T^{\prime}_{j,:,:}\end{bmatrix}_{0\leq j<J}=U^{\intercal}\begin{bmatrix}B_{r,:}\times C_{r,:}\end{bmatrix}_{r}$.
The upshot of constructing $T^{\prime}$ is that if there exists $A,B,C$ s.t. $T=\sum_{r}A_{r}\times B_{r,:}\times C_{r,:}$, then $\forall i,\ T_{i,:,:}\in\textrm{span}\left\{B_{r,:}\times C_{r,:}\right\}_{0\leq r<R}$, so $\left\{T_{i,:,:}\right\}_{0\leq i<n}$ spans a linear subspace of $\mathbb{F}^{n\times n}$ with rank at most $R$. This forces $T^{\prime}$ to have at most $R$ many slices, i.e., $J\leq R$, so there are at most $|\mathbb{F}|^{R^{2}}$ possible $U$ to fix. Thus, we avoid exponential dependence on $n$.
If $J=0$, then $T$ must be all-zero, so it trivially has a rank-$R$ decomposition (e.g., with all terms set to 0). Otherwise, for every possible $U$, run Gaussian elimination on $\Xi:=\left[\begin{array}[]{c|c}U^{\intercal}&\begin{bmatrix}\vdots\\ T^{\prime}_{j,:,:}\\ \vdots\end{bmatrix}_{0\leq j<J}\end{array}\right]$ to obtain
$\Xi^{\prime}:=\left[\begin{array}[]{c|c}\Omega&\begin{bmatrix}\vdots\\ D_{j}\\ \vdots\end{bmatrix}_{0\leq j<J}\end{array}\right]$, for $\Omega:=\mathrm{rref}\left(U^{\intercal}\right)$ and some $n\times n$ matrices $D_{j}$.
If there is some row $r$ s.t. $\Omega_{r,:}$ is all-zeros but $D_{r}$ is not, then there is no solution. Otherwise, since all all-zero rows of a rref matrix must be at the bottom, we can remove everything except the first $N$ rows of $\Xi^{\prime}$, where $N$ is the number of nonzero rows of $\Omega$.
Now we analyze all possible forms of the truncated rref matrix $E:=\Omega_{:N,:}$ with $R$ columns and $N\leq J\leq R$ rows, for $R\leq 3$:
  • For $R=1$, we must have $E=\begin{bmatrix}x\end{bmatrix}$, so we simply check whether $\mathrm{rank}\left(D_{0}\right)\leq\begin{cases}1&x\neq 0\\ 0&\textrm{else}\end{cases}$.
  • For $R=2$, we either have $E=\begin{bmatrix}1\\ &1\end{bmatrix}$, in which case we check whether $\mathrm{rank}\left(D_{0}\right)\leq 1\ \land\ \mathrm{rank}\left(D_{1}\right)\leq 1$; or $E=\begin{bmatrix}1&x\end{bmatrix},\begin{bmatrix}0&1\end{bmatrix}$, in which case we check whether $\mathrm{rank}\left(D_{0}\right)\leq(\#\textrm{ of nonzeros in }E_{0,:})$.
  • For $R=3$, there are a few more cases, of which one is nontrivial:
    • $E=\begin{bmatrix}1&x&y\end{bmatrix},\begin{bmatrix}0&1&x\end{bmatrix},\begin{bmatrix}0&0&1\end{bmatrix}$: check whether $\mathrm{rank}\left(D_{0}\right)\leq(\#\textrm{ of nonzeros in }E_{0,:})$;
    • $E=\begin{bmatrix}1\\ &1\\ &&1\end{bmatrix}$: check whether $\mathrm{rank}\left(D_{i}\right)\leq 1\ \forall i=0,1,2$;
    • $E=\begin{bmatrix}1&x\\ &&1\end{bmatrix},\begin{bmatrix}0&1\\ &&1\end{bmatrix}$: check whether $\mathrm{rank}\left(D_{i}\right)\leq(\#\textrm{ of nonzeros in }E_{i,:})\ \forall i=0,1$;
    • $E=\begin{bmatrix}1&&x\\ &1&y\end{bmatrix}$ (the nontrivial case):
      If $x=0$ or $y=0$, we can do the same thing as in $\begin{bmatrix}1&x\\ &&1\end{bmatrix}$ or $\begin{bmatrix}0&1\\ &&1\end{bmatrix}$, as the sets of nonzero terms of different rows in $E$ are pairwise disjoint.
      Otherwise, check whether $\exists i\textrm{ s.t. }\mathrm{rank}\left(D_{i}\right)=2$ (notice the $=$ instead of $\leq$ this time):
        if such $i$ exists, enumerate all rank-2 decompositions of $D_{i}=u_{0}\times v_{0}+u_{1}\times v_{1}$, and for each decomposition, check whether $\mathrm{rank}\left(D_{j}-u_{1}\times v_{1}\right)\leq 1$ for the other index $j\neq i$;
        otherwise, $\mathrm{rank}\left(D_{0}\right)\leq 1$ and $\mathrm{rank}\left(D_{1}\right)\leq 1$, so we can force $B_{2,:},C_{2,:}=\overrightarrow{0}$ and solve for $D_{0}=B_{0,:}\times C_{0,:},\ D_{1}=B_{1,:}\times C_{1,:}$;
In contrast to general tensors, rank decomposition of a $n\times n$ matrix (with the decomposition having smallest possible rank) can be computed in $O(n^{3})$ time via Gaussian elimination, so if all relevant checks in this casework pass, it is easy to construct a solution for $B,C$.
It remains to count how many rank-2 decompositions there are for a rank-2 matrix. Fortunately we have such a bound for arbitrary matrix rank which does not have exponential dependence on $n$:
Theorem 3.
Consider an $n\times m$ matrix $M$ over an arbitrary field with rank $r:=\mathrm{rank}\left(M\right)>0$. Let $F_{0}=\mathrm{rref}\left(M\right)_{:r,:}$ and $C_{0}$ satisfy $C_{0}F_{0}=M$. Then for any $n\times r$ matrix $C$ and $r\times m$ matrix $F$ s.t. $CF=M$, there is some invertible $r\times r$ $X$ s.t. $C=C_{0}X^{-1}$ and $F=XF_{0}$.
@@
First we note that $C_{0}$ exists and can be constructed via Gaussian elimination by maintaining the composition of all row operations in a separate matrix (and then truncating that matrix to its first $r$ columns). Additionally, $C_{0}$ is unique, because $\mathrm{rank}\left(F_{0}\right)=r$ and $r\leq\min\left(n,m\right)$ force the rows of $F_{0}$ to be linearly independent.
We have $\mathrm{rowspan}\left(M\right)=\mathrm{rowspan}\left(CF\right)\subseteq\mathrm{rowspan}\left(F\right)$ and $\mathrm{rank}\left(F\right)\leq r$, so $\mathrm{rowspan}\left(M\right)=\mathrm{rowspan}\left(F\right)$. Since row reduction and removing all-zero rows preserves row-span, $\mathrm{rowspan}\left(F_{0}\right)=\mathrm{rowspan}\left(\mathrm{rref}\left(F\right)\right)$.
We recall a property of rref matrices. Let $W$ be an arbitrary rref matrix with no all-zero rows. By definition of rref, all terms directly above and below any leading 1 must be 0, so in any linear combination of rows of $W$, leading 1s cannot be canceled out with any other terms. Thus, denoting $l_{i}$ as the column position of the leading 1 at row $i$, the position of the leftmost nonzero element of the row vector $\overrightarrow{v}:=\sum_{i}c_{i}W_{i,:}$ for scalars $c_{i}$ is $l_{k}$ for $k=\min{\left\{i:c_{i}\neq 0\right\}}$. Furthermore, $\overrightarrow{v}_{l_{i}}=c_{i}$, so given an arbitrary linear combination $\overrightarrow{v}$ of the rows of $W$ it is possible to determine the sequence of coefficients $c_{i}$ used to obtain the linear combination. Since the rows of $W$ are linearly independent, there is exactly one such sequence of coefficients $c_{i}$.

### v3

@@ Abstract
We show that finding rank-$R$ decompositions of a 3D tensor, for $R\leq 4$, over a fixed finite field can be done in polynomial time. However, if some cells in the tensor are allowed to have arbitrary values, then rank-2 is NP-hard over the integers modulo 2. We also explore rank-1 decomposition of a 3D tensor and of a matrix where some cells are allowed to have arbitrary values.
@@ 1.1 Results
For $R\leq 4$, finding a rank-$R$ decomposition over a finite field $\mathbb{F}$ of an $n\times n\times n$ tensor, or determining that it does not exist, can be done in $O(f(|\mathbb{F}|,R)n^{3})$ time, for some function $f$.
We show tighter runtime analysis in Section 2.3.1.
@@
1.2 Notation
  • We use NumPy notation for indexing on tensors.
    • e.g. for a matrix $M$:
        $M_{i,:}$ denotes the $i$-th row
        $M_{:i,:}$ denotes the matrix truncated to its first $i$ rows
        $M_{:,i}$ denotes the $i$-th column
  • All tensors and sequences are 0-indexed.
  • The reduced row echelon form of a matrix $M$ is denoted as $\mathrm{rref}\left(M\right)$.
@@ 2 Without wildcards
2.1 Preliminary results about matrix factorizations
Before describing our algorithm for Theorem 1, we present several lemmas on matrix rank decompositions that we will use extensively.
The first lemma is a direct result of Gaussian elimination:
Lemma 1.
Given any matrix $M\in\mathbb{F}^{m\times n}$, $\exists T\in\mathrm{GL}\left(m,\ \mathbb{F}\right)$ s.t. $TM=\mathrm{rref}\left(M\right)$.
It follows that inverting a $n\times n$ matrix can be done in $O(n^{3})$ time.
The next lemma is the uniqueness of rref (up to all-zero rows):
Lemma 2.
Two rref matrices with the same rowspan must be equal, up to the presence of all-zero rows.
@@ 2.1 Preliminary results about matrix factorizations
Consider an rref matrix $W=\left[\begin{array}[]{c}w_{0}\\ \hline\cr\vdots\\ \hline\cr w_{k-1}\\ \hline\cr O\end{array}\right]$, where each $w_{i}\neq\overrightarrow{0}$ has its leading 1 at column $\ell_{i}$, with $\ell_{0}<\dots<\ell_{k-1}$. Consider a linear combination of the rows, $v:=\sum_{i}\alpha_{i}w_{i}$, where the $\alpha_{i}$ are not all zero: then $v_{\ell_{i}}=\alpha_{i}$, since by definition of rref, every column of $W$ that has a leading 1 has all of its other elements set to 0. As a consequence, if we are instead given some $v$ that is known to be a linear combination of $w_{i}$, then the list of coefficients $\alpha_{i}$ of that linear combination are unique and can immediately be deduced by reading each $v_{\ell_{i}}$ element.
Additionally, the leading nonzero term of $v$ must be at some $\ell_{i}$-th column, as we can use the following casework: if $\alpha_{0}\neq 0$, the leading nonzero term of $v$ is $\ell_{0}$, since each $w_{i}$ is all zeros to the left of its leading 1: else if $\alpha_{1}\neq 0$, the leading nonzero term of $v$ is $\ell_{1}$: etc.